OpenStudy (anonymous):

Find the equation of the tangent line and normal line to the curve y= (6+3x)^2 at the point (1,81). Tangent line: y= Normal line: y=

OpenStudy (anonymous):

y'=2(6+3x)*3=18x+36 y'(1)=18+36=54=slope

OpenStudy (anonymous):

m=(y-y_1)/(x-x_1) 54=(y-81)/(x-1) find y..

OpenStudy (anonymous):

Do you need more help?

OpenStudy (anonymous):

Yes Please!

OpenStudy (anonymous):

Do you understand the slope value 54?

OpenStudy (anonymous):

Yes (:

OpenStudy (anonymous):

When you want to find tangent => find first derivative!

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

Value of slope => constant number of derivative at tangent point!

OpenStudy (anonymous):

Do you follow me?

OpenStudy (anonymous):

I do yes!

OpenStudy (anonymous):

Can you find derivative for me?

OpenStudy (anonymous):

18(x+2)

OpenStudy (anonymous):

Tell me where you have the trouble?

OpenStudy (anonymous):

Yep, you're right :) f'(x) = 18 x + 36

OpenStudy (anonymous):

Since tangent line is linear which y = mx + b => We need slope m, and the tangent point!

OpenStudy (anonymous):

Can you find slop m?

OpenStudy (anonymous):

Not sure how to

OpenStudy (anonymous):

Just plug the value of tangent point in: At ( 1, 81), f'(1) = ...

OpenStudy (anonymous):

54

OpenStudy (anonymous):

Yep f'(1) = 18 * 1 + 36 = 54 Are you sure you understand how to find slope now?

OpenStudy (anonymous):

Yes thank you for your help